Soybean meal is 16% protein;cornmeal is 8% protein .How many pounds of each should be mix together in order to get 320-lbs mixturethat is 13% protein.
How many pounds of soybean should be in the mixture._____
How many pounds of cornmeal should be in the mixture._____


Answer by checkley77(12844)

.16x+.08(320-x)=.13*320
.16x+25.6-.08x=41.6
.08x=41.6-25.6
.08x=16
x=16/.08
x=200 lbs. of soybeans is used.
320-200=120 lbs. of cornmeal is used.Proof:
.16*200+.08*120=41.6
32+9.6=41.6
41.6=41.6